OpenStudy (pphalke):

h=17, b=9 and B=11 a=1/2(b+B)h find the value of A given that

OpenStudy (kayders1997):

Plug in the given values

hartnn (hartnn):

so where are you stuck at?

hartnn (hartnn):

9 + 11 ? or the division by 2?

OpenStudy (mathmale):

Your post, " 54 pphalke h=17, b=9 and B=11 a=1/2(b+B)h find the value of A given that would be a lot clearer if you could provide any illustration that came with these equations. What do all these variables represent? h, b, B, a, h, A

Directrix (directrix):

A=1/2(b+B)*h appears to be a formula for finding the area of a trapezoid. A=1/2*(9+11)*17 @pphalke

OpenStudy (pphalke):

i did that nbut now what

Directrix (directrix):

9+11= ?

Directrix (directrix):

@pphalke ^^^

OpenStudy (pphalke):

20 @direct x

Directrix (directrix):

Good. 1/2 of 20 = ?

OpenStudy (pphalke):

I don't know

Directrix (directrix):

Type this in the Google search box and click return. The Google calculator will pop up with the answer. 1/2 of 20 =

OpenStudy (pphalke):

10

OpenStudy (pphalke):

@directx?

OpenStudy (pphalke):

@Directrix does that mean 10 is the answer?

Directrix (directrix):

Almost. One more calculation to get it.

Directrix (directrix):

10 *17 =

OpenStudy (pphalke):

whats that

OpenStudy (pphalke):

170

Directrix (directrix):

10 times 17 =

OpenStudy (pphalke):

its 170 10*17

Directrix (directrix):

I put the formula parts back together so that you can see the total formula worked out. A=1/2*(9+11)*17 A = 1/2 * (20)*17 A = 10 * 17 A = 170 -->

OpenStudy (pphalke):

is that the answer

Directrix (directrix):

Correct

Directrix (directrix):

Question about it?